OpenStudy (anonymous):

Find the derivative. f(x)= tan(3x)(x^2+1)

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad (x^2+1) \tan(3x)\]Derivative? Ok we have the `product` of two x-thingies. Product indicates we should apply a certain rule, yes? Hmmmmm

OpenStudy (anonymous):

So you'd use the (f'(g(x))(g'(x))?

zepdrix (zepdrix):

That's called `chain rule`. Yes we will eventually need to use that. But that's not what we want to start with. Our function consists of a product of functions of x. So we need to start by applying the `product rule`.\[\Large\bf\sf f'(x)\quad=\quad \color{royalblue}{\left[x^2+1\right]'}\tan(3x)+(x^2+1)\color{royalblue}{\left[\tan(3x)\right]'}\]

zepdrix (zepdrix):

That's our setup for product rule. We need to take the derivative of the blue parts.

OpenStudy (anonymous):

So the derivative would be 2x for the first one and 3sec^2x?

zepdrix (zepdrix):

Oh oh one small boo boo..

OpenStudy (usukidoll):

*eats zepdrix's head and lets go*

zepdrix (zepdrix):

\[\Large\bf\sf f'(x)\quad=\quad \color{orangered}{\left[2x\right]}\tan(3x)+(x^2+1)\color{orangered}{\left[3\sec^2(\color{royalblue}{x})\right]}\]

zepdrix (zepdrix):

We need to fix that part in blue.

zepdrix (zepdrix):

Ahh my head D:

OpenStudy (usukidoll):

missing 3x

zepdrix (zepdrix):

When you take the derivative of the `outer function` (tangent), the inner function should NOT change. right?

OpenStudy (usukidoll):